The stock of Apple is currently selling for $55/share. If the 3 month risk-free rate is 5% p.a. semi-annual compounding, what is the forward price of a 3-month forward contract to buy one share of Apple?
Answer #1
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+5/(2*100))^2-1)*100
Effective Annual Rate% = 5.0625
Future value = present value*(1+ rate)^time
Future value = 55*(1+0.050625)^0.25
Future value = 55.6833
Using Calculator: press buttons "2ND"+"FV" then assign
PV =-55
I/Y =5.0625
N =0.25
PMT = 0
CPT FV
Using Excel
=FV(rate,nper,pmt,pv,type)
=FV(0.050625,0.25,,-55,)